A gaseous nitrogen sample in a 78 L air bag has a pressure of 800 mmHg. If the sample is transferred into a 30 L bag at same temperature, what will be the value of its new pressure? 


1
Expert's answer
2021-02-01T04:05:26-0500

pV=nRT. n= pV/RT =78*800/62400*300= 0,00333333333.

p= nRT/V = 0,00333333333 x 62400 x 300/30 = 2 080 mmhg